OpenStudy (anonymous):

Solve the exponential equation. Express your solutions in exact form only. Please show all of your work. 14e^x log9 =5 Solve the exponential equation. Express your solutions in exact form only. Please show all of your work. 14e^x log9 =5 @Mathematics

OpenStudy (anonymous):

\[14e^xlog9=5\]

OpenStudy (anonymous):

5/(14*log9) = e^x x= log(5/(14*log(9)) x= -1.817

OpenStudy (anonymous):

Thankyou. Is this the solution?